# spacing=nonfrench

### v2

@@ 7.5 Comparison with GPUs
MeshGEMV(WSE-2) Time (ms) 0.0012 0.00203
cuBLAS(A100) Time (ms) 0.336 1.231
WSE-2/A100 Energy Ratio 10.37 22.46
Table 6: Comparing MeshGEMV(WSE-2) with cuBLAS(A100) in GEMV latency and energy.
GEMM [16K,16K]$\times$[16K,16K] [32K,32K]$\times$[32K,32K]
MeshGEMM(WSE-2) Time (ms) 4.8 34
cuBLAS(A100) Time (ms) 34.4 282.1
WSE-2/A100 Energy Ratio 0.265 0.307
Table 7: Comparing MeshGEMM(WSE-2) with cuBLAS(A100) in GEMM latency and energy.
Finally, we compare WaferLLM against the SOTA LLM inference system on GPUs. For this experiment, we use Cerebras WSE-2 and NVIDIA A100, both manufactured on TSMC’s 7nm process, ensuring a fair comparison. To compare against the H100 fairly, we would need access to the WSE-3 which is unavailable for us.
GEMV. We compare MeshGEMV with GEMV on GPUs, isolating the difference in numerical operators that Cerebras has yet to optimize fully compared to CUDA. Shown by Table 6, for a cuBLAS implementation GEMV [32], MeshGEMV outperforms GPU by 606$\times$ in completion time, showcasing the advantages of providing substantial memory bandwidth through wafer-scale devices. This translates to 22$\times$ greater energy efficiency, reflecting the benefits of wafer-based connections (connecting on-chip memory) over PCB-based ones (connecting off-chip HBM) in GPUs.
@@
GEMM. Cerebras WSE-2 offer limited benefits for GEMM, since energy-efficient wafer-based connections do not address its compute bottleneck. Performance ultimately hinges on the density of transistors and the efficiency of the core design, which is reflected by the specification of Cerebras WSE-2 (i.e., 50% less energy efficient than A100 GPU).
In absolute performance, MeshGEMM is 8.3$\times$ faster than CUDA GEMM, as shown in Table 7, largely due to its larger on-chip area. However, this speedup does not translate into energy savings. MeshGEMM is observed to be 70% less energy efficient than GPU GEMM, a gap expected to narrow with improved core designs, as seen in Cerebras WSE-3, which increases core efficiency by 100%.
Sequence Input 4096 / Output 4096 LLaMA3-8B LLaMA2-13B
WaferLLM(WSE-2) Throughput (tokens/s) 2480.4 1848
vLLM(A100) Throughput (tokens/s) 78.36 47.86
WSE-2/A100 Energy Ratio 1.41 1.71
Table 8: Comparing WaferLLM(WSE-2) with vLLM(A100) in end-to-end throughput and energy.
LLM inference. Finally, we compare WaferLLM with vLLM for running LLM inference, shown as Table 8. Unlike isolated GEMV and GEMM comparisons, full LLM inference include numerous system components that WaferLLM does not optimize and are limited by the current maturity of the Cerebras software stack (e.g., less optimized numerical operators compared to NVIDIA CUDA). Despite these constraints, WaferLLM achieves 2480 tokens/s and 1848 tokens/s throughput for LLaMA3-8B and LLaMA2-13B, respectively, translating to 31.6$\times$ and 38.6$\times$ speedup and approximately 1.4$\times$ and 1.7$\times$ better energy efficiency in tokens per joule.
The reduction in energy efficiency from GEMV’s 22$\times$ to LLM’s 1.7$\times$ stems from two factors: (i) WSE-2 cores have limited local SRAM (48KB), preventing efficient tensor parallelism as in vLLM and necessitating pipeline parallelism, which introduces execution bubbles and reduces chip utilization by 5$\times$; and (ii) LLaMA models are optimized for GPU architectures, with narrow layers designed to minimize off-chip overhead. This limits layer placement across WSE-2 cores, exacerbating bubble issues.
8 Current Limitations and Future Directions
@@
Execution bubbles. The performance of WaferLLM is currently constrained by execution bubbles caused by the need for pipeline parallelism. Increasing a core’s compute efficiency and local memory could mitigate the need for pipeline parallelism, potentially allowing the wafer-scale chip to decode over 10,000 tokens per second with LLMs.
LLM model design. The limited bandwidth of HBM has constrained today’s LLM model designs from using large tensors in transformer layers[16, 5]. With the advent of wafer-scale accelerators, we anticipate new LLM architectures adopting larger tensors (for example, the architectures with significantly wider layers than those today), free from hardware limitations, and avoiding the bubble issues when supporting current LLMs in WaferLLM.
Beyond Cerebras WSE. Although evaluated with Cerebras WSE, the PLMR model applies to future devices such as Tesla Dojo, featuring hundreds of thousands of cores connected via a mesh, each with MBs of local memory and limited NoC routing. Future devices may adopt different mesh-like architectures with shorter NoC paths, such as 2D torus or hybrids with on-chip switches, all consistent with the PLMR model. Our wafer-scale LLM parallelism will continue to support these architectures effectively. MeshGEMM and MeshGEMV address the worst-case 2D mesh scenario and will remain better, at least not worse, than baseline methods.
TSMC’s newer System-on-Wafer integration, expected in 2027, could boost chip density by 40$\times$ on a wafer while still aligning with the PLMR model, ensuring the long-term relevance of our contributions.
@@ 9 Related Work
Wafer-scale allreduce. Recent research [26] has investigated wafer-scale allreduce, but a single allreduce cannot fully parallelize GEMV or support full LLM inference as achieved by WaferLLM. Additionally, this prior work is a specific instance of the two-way K-tree allreduce proposed in WaferLLM.

### v3

@@ 7.5 Comparison with GPUs
TP in SGLang(A100) Time (ms) 1 GPU 8 GPUs 2$\times$8 GPUs 1 GPU 8 GPUs 2$\times$8 GPUs
0.336 0.253 0.340 1.231 0.341 0.339
MeshGEMV (WSE-2)
Time (ms)
0.0012 0.00203
A100/WSE-2 Energy Ratio
7.47 44.97 120.88 16.17 35.83 71.25
Table 6: Comparing MeshGEMV (WSE-2) with TP in SGLang(A100) of GEMV latency and energy.
We compare WaferLLM against the state-of-the-art LLM inference system on GPUs using Cerebras WSE-2 and NVIDIA A100. To compare against the H100 fairly, we would need access to the WSE-3, which is a 5nm manufacturer but unavailable to us.
GEMV. We compared MeshGEMV with the GEMV parallelization strategy that follows SGLang’s [53] multi-GPU tensor parallelism, while the computation on each GPU is accelerated using the cuBLAS library. Shown by Table 6, compared to a single A100 GPU, MeshGEMV outperforms by 280-606$\times$ with different matrix sizes, showcasing the advantages of providing substantial memory bandwidth through wafer-scale devices. This also translates to 7.5-16$\times$ greater energy efficiency, reflecting the benefits of wafer-based connections (connecting on-chip memory) over PCB-based ones (connecting off-chip HBM) in GPUs.
Distributed GEMV on multi-GPUs shows limited scalability: a single A100 yields the highest energy efficiency per FLOP, and performance improves only 1.32$\times$ from one to eight GPUs, then degrades at sixteen. This inefficiency stems from memory intensity and communication overhead over NVLink and IB. In contrast, MeshGEMV exploits WSE-2’s low-latency NoC to scale across hundreds of thousands of cores, achieving 166-210$\times$ higher performance and 45-70$\times$ better energy efficiency at peak compared to the best GPU cluster results. Moreover, GEMV exhibits similar performance for 16K and 32K tasks on two machines with 16 GPUs, mainly because communication overhead dominates the computation in distributed GEMV across two nodes.
@@
Prefill (4K CTX) LLaMA3-8B LLaMA2-13B
SGLang(A100) TPR 1 GPU 8 GPUs 2$\times$8 GPUs 1 GPU 8 GPUs
13988 17361 13994 7805 12287
WaferLLM WSE-2 TPR
27686 17498
A100/WSE-2 Energy Ratio 0.05 0.34 0.84 0.06 0.30
Table 7: Comparing WaferLLM (WSE-2) with SGLang(A100) in prefill throughput and energy.
Decode (4K CTX) LLaMA3-8B LLaMA2-13B
SGLang(A100) TPR 1 GPU 8 GPUs 2$\times$8 GPUs 1 GPU 8 GPUs
78 260 164 48 175
WaferLLM WSE-2 TPR
2700 2039
A100/WSE-2 Energy Ratio 0.92 2.22 7.02 1.13 2.49
Table 8: Comparing WaferLLM (WSE-2) with SGLang(A100) in decode throughput and energy.
LLM inference. The maximum TPR achievable on GPU clusters is significantly lower than on WSE-2. We compare WaferLLM with SGLang on an A100 multi-GPUs cluster and find that WaferLLM delivers a 6-20$\times$ higher TPR across a range of input/output lengths and model sizes from 8B to 72B, with the gap widening for longer outputs and larger models (Table 2). As shown in Tables 7 and 8, scaling SGLang from 1 to 8 GPUs yields only 1.2-1.6$\times$ prefill and 3.3-3.6$\times$ decode speedups, far below ideal linear scaling. Performance further degrades when scaling to 16 GPUs, due to inter-node communication bottlenecks. As a result, peak TPR for dense models is typically reached within a single 8-GPU node, where WaferLLM still outperforms SGLang up to 20$\times$. For use cases demanding high per-request throughput, wafer-scale accelerators are substantially more capable than conventional xPU-based systems.
In terms of energy efficiency, though a WSE-2 chip has  47$\times$ the area and  37$\times$ the power and cost of an A100, WaferLLM still achieves a 2-2.5$\times$ energy efficiency advantage at SGLang’s optimal multi-GPU result, owing to GPUs’ nonlinear scaling on decode. This advantage is especially valuable for long-output scenarios like test-time scaling, where serving cost is critical. However, compared to GEMV, WaferLLM’s decode energy efficiency advantage is reduced due to: (i) limited local SRAM (48KB) on WSE-2 cores, which hinders efficient tensor parallelism and necessitates pipeline parallelism, causing up to 5$\times$ underutilization; and (ii) GPU-optimized LLaMA models with narrow layers that constrain layer placement and worsen utilization on WSE-2.
8 Implementation Detail and Future Direction
@@
Hardware architecture. The performance of WaferLLM is currently constrained by execution bubbles caused by the need for pipeline parallelism. Increasing a core’s local memory by 5-6$\times$ could eliminate the need for pipeline parallelism, enabling full tensor parallelism, as on vLLM and SGLang. Wafer-scale chip designers are already moving in this direction. Cerebras WSE-3 retains the same NoC configuration but improves per-core efficiency and local memory, while Tesla’s Dojo incorporates 1MB of per-core memory.
Memory-to-Compute Ratio. LLM decoding demands a near 1:1 memory-to-compute ratio. However, GPUs like A100 have limited on-chip SRAM, forcing frequent off-chip memory access and yielding a poor ratio of 1:312 (FP16). Multi-GPU setups exacerbate this due to heavy inter-GPU communication. In contrast, Cerebras WSE-2 maps most model weights onto on-chip memory via a low-latency NoC, achieving near-ideal locality and approaching a 1:1 ratio. To fully realize this balance on mesh-based NoCs, adherence to the PLMR model is essential, enabling WaferLLM to outperform GPU-based systems by orders of magnitude in TPR.
Handle reliability issues. Currently, Cerebras WSE-2/3 handles faults by hardware, only exposing healthy cores (organized in a mesh) to software, with no explicit handling required at the software level. Moreover, redundant cores and links are built in at fabrication, and the on-chip SoC dynamically remaps and reroutes around defects at runtime, ensuring minimal performance impact at a low redundancy cost. Meanwhile, wafer-scale chip makers recently reported a 93% functional wafer area, higher than 70–80% in commercial GPUs [8], due to the smaller area per core design. Over two years of WSE-2 deployment, we have observed high reliability, confirming the effectiveness of these fault-tolerance mechanisms in real-world use.
Various model architecture. WaferLLM is also beneficial for MoE as it shares key operators with dense LLMs, including MeshGEMM, MeshGEMV, and shift-based KV cache management. The main difference is the all-to-all communication between attention and expert layers, which we implement using WSE-2’s NoC multi-cast operations. Further optimizations for sparse models, such as offloading and sparse attention, are among our future research.
Beyond Cerebras WSE. While evaluated on Cerebras WSE, the PLMR model generalizes to emerging mesh-like architectures such as Tesla Dojo, which also feature hundreds of thousands of cores with local memory and constrained NoC routing. Variants like 2D torus or hybrid mesh-switch topologies also conform to PLMR. Our design for MeshGEMM and MeshGEMV targets worst-case 2D mesh and remains competitive across such platforms. Beyond on-chip meshes, chip-to-chip mesh interconnects, as seen in Tenstorrent’s core- and card-level meshes, also align well with PLMR. Looking ahead, advancements in wafer-scale integration, such as TSMC’s projected 40$\times$ density increase by 2027, further reinforce the long-term applicability of our approach.
@@ 9 Related Work
Wafer-scale allreduce. Recent research [27] has investigated wafer-scale allreduce, but a single allreduce cannot fully parallelize GEMV or support full LLM inference as achieved by WaferLLM. Additionally, this prior work is a specific instance of the K-tree allreduce proposed in WaferLLM.
@@ 10 Conclusion
11 Acknowledgments
We sincerely thank our shepherd, Marco Canini, and the OSDI reviewers for their insightful feedback that greatly improved this paper. We are also grateful to Yuqing Xia (Microsoft) for her valuable input during the development of this work.
We acknowledge the hardware resources provided by the Edinburgh International Data Facility (EIDF) and the Edinburgh Parallel Computing Centre (EPCC), particularly their support in granting access to the Cerebras WSE systems. We thank Nick Johnson for coordinating access and Mark Parsons for supporting the hardware acquisition and our project.
Finally, we appreciate the generous assistance of Leighton Wilson and Mathias Jacquelin from Cerebras for their patience and detailed responses to our many questions regarding CSL and the Cerebras WSE.